Find three consecutive integers whose sum is -33.
Found 2 solutions by Alan3354, nerdybill:
Answer by Alan3354(69443)

-33/3 = -11, the middle number

Answer by nerdybill(7384)

Find three consecutive integers whose sum is -33.
.
Let x = 1st integer
then
x+1 = 2nd integer
x+2 = 3rd integer
.
x + x+1 + x+2 = -33
3x + 3 = -33
3x = -36
x = -12 (first integer)
x+1 = -11 (2nd integer)
x+2 = -10 (3rd integer)
.
The three integers are: -12, -11 and -10